OpenStudy (anonymous):

What is the sum of a 17-term arithmetic sequence where the first term is 8 and the last term is -88? -542 -680 -893 -954

OpenStudy (ddcamp):

Sum of an arithmetic sequence = \[\frac{ 1 }{ 2 }(Number\ of\ terms)(First\ term+Last\ term)\] \[=\frac{ 1 }{ 2 }(17)(-88 + 8) = 8.5(-80) = -680\]